OpenStudy (anonymous):

Linear algebra question: a vector u and a set S are given. If possible, write u as a linear combination of the vectors in S. u = [3], S = {[1],[2],[-2]} [8], {[2],[3],[-5]} Thank you.

OpenStudy (amistre64):

1 2 -2 |3 2 3 -5 |8 and augment

OpenStudy (amistre64):

Sx = u x = aug(S|u)

OpenStudy (amistre64):

1 2 -2 |3 2 3 -5 |8 1 2 -2 | 3 -1 -3/2 5/2 |-4 ---------------- 0 1/2 1/2 -1 1 2 -2 | 3 0 1 1 |-2 1 0 -4 | 7 0 1 1 |-2 maybe x1 4 7 x = x2 = x3 -1 + -2 x3 1 0

OpenStudy (amistre64):

with x3 = 0; we can use: Sx = 7s1 -2s2 + 0s3 = u

OpenStudy (anonymous):

I'm a bit confused. How did you figure out to set x3=0?

OpenStudy (amistre64):

becasue x3 can be any value; so I simply choose the easiest one for me to use

OpenStudy (amistre64):

the row reduction also tells me that I only really need to use the first 2 column vectors so the last ones a but of a useless tagalong

OpenStudy (anonymous):

Got it. Thank you.